# Neural Wave Functions for Superfluids

### v3

@@ VI Discussion
In this work, we studied the benchmark superfluid system of the UFG with neural wave functions 555 Shortly after submitting this manuscript to the arXiv server, a closely related preprint on neural networks and the Fermi gas appeared [68], proposing a similar extension to continuous-space neural network ansatzes. . We showed that the original FermiNet Ansatz has difficulties in describing such a paired systems with strong and short-range attractive interactions between opposite spin particles. Hence, we proposed a way to improve the variational Ansatz by using determinants of FermiNet geminals, similar to an AGPs or a BCS wave function. We showed mathematically that the original FermiNet is a limiting case of the AGPs FermiNet despite the use of fewer parameters in the latter, which means that any FermiNet wave function can in-principle be written as an AGPs FermiNet wave function.
@@
We compared the total energies and the energies per particle of the UFG between the original FermiNet and the AGPs FermiNet. The former fails to produce a paired state when the number of particles, $N$, is greater than around $10$, while the AGPs FermiNet works very well.
@@
As the UFG has a superfluid ground-state, we computed the pairing gap and condensate fraction for the $N=38$ system and compared estimates made with the original FermiNet and the AGPs FermiNet. There is a clear qualitative difference between the pairing gap obtained using the AGPs FermiNet and the original FermiNet, with the latter approaching zero as the number of particle $N$ increases. The results for the superfluid condensate fraction show a similar behaviour, with the AGPs FermiNet converging to a finite value, while the original FermiNet approaches zero in the limit of large system size. Although the AGPs pairing gap shows significant finite size errors, it lies close to the mean-field BCS result with Gorkov-Melik-Barkhudarov corrections [57]. Taken together, these results show that the original FermiNet is unable to represent large systems with superfluid ground states. The AGPs FermiNet is much more suitable for studying paired systems such as the UFG.
@@
To demonstrate success of the AGPs FermiNet, we also compared our calculated total energies with state-of-the-art fixed-node diffusion QMC energies obtained using a Jastrow-BCS Ansatz [21]. For all systems with more than a few particles, the AGPs FermiNet achieves lower (i.e., better) variational energies than FN-DMC using the same model interaction and system parameters.
@@
The failure of the original FermiNet Ansatz comes as a surprise because the original FermiNet paper [1] argued that any many-body fermionic wave function can be represented as a single determinant of FermiNet orbitals. However, the mathematical argument relies on the construction of FermiNet orbitals with unphysical discontinuities. Whether or not any wave function can be represented as a single determinant of FermiNet orbitals of the type used in practice, which are differentiable everywhere except at electron-electron and electron-nuclear coalescence points, remains an open question. Another limitation is that the original FermiNet architecture, which is rather simple, may not be able to represent an arbitrary FermiNet orbital.
Even if a single FermiNet determinant is general in principle, there is no guarantee that it is equally easy to represent all wave functions. It may be that producing an accurate representation of a paired wave function requires the width and number of layers in the network to increase rapidly with system size. The observation that the original FermiNet works well when $N\lessapprox 10$ but that the quality of the results degrades rapidly for larger systems suggests that this is, in fact, the case.
One of the most important strengths of the original FermiNet is that it does not require the use of an explicit basis set. In particular, there is no need to construct and optimize a new basis set for every new system or particle. Our modified AGPs FermiNet for paired superfluids inherits this advantage.
@@
Another strength of the AGPs FermiNet is the ease with which it is possible to optimize linear combinations of determinants of FermiNet pairing orbitals, such as the one in Eq. (43). This is much more difficult to accomplish with conventional wave functions based on explicit two-electron pairing orbitals or pairing orbitals represented as outer products of single-particle orbitals or basis functions. The basis-set-free nature of the AGPs FermiNet will make it relatively easy to investigate the importance of pairing in other systems of interest such as molecules, electron-positron systems, electron-hole liquids, and other $s$-wave superfluids.
The AGPs FermiNet introduced here is very general and flexible. It can be generalized to arbitrary pairing also with triplet spin components. Therefore, we expect it to become a powerfull tool for understanding strongly correlated non-$s$-wave superfluid and superconducting systems as well, for example Helium-3 or high-$T_{c}$ and $p$-wave superconductors.
Besides the FermiNet, future work may involve utilizing more recent neural-network architectures, such as the Psiformer [65], GLOBE and MOON [66], DeepErwin [67], etc., to build neural-network-based AGPs or Pfaffian wave functions.
@@ A.1 Fixed-particle Number BCS Ground State
|$\ket{\Psi_{\text{BCS}}}$ $=\prod_{\mathbf{k}}(u_{\mathbf{k}}+v_{\mathbf{k}}\hat{c}^{\dagger}_{\mathbf{k},\uparrow}\hat{c}^{\dagger}_{-\mathbf{k},\downarrow})\ket{0}$
$=\bigg{(}\prod_{\mathbf{k}}u_{\mathbf{k}}\bigg{)}e^{\sum_{\mathbf{k}}\varphi_{\mathbf{k}}\hat{c}^{\dagger}_{\mathbf{k},\uparrow}\hat{c}^{\dagger}_{-\mathbf{k},\downarrow}}\ket{0},$
@@
|$\ket{\Psi_{\text{PBCS}}}=\bigg{(}\sum_{\mathbf{k}}\varphi_{\mathbf{k}}\hat{c}^{\dagger}_{\mathbf{k},\uparrow}\hat{c}^{\dagger}_{-\mathbf{k},\downarrow}\bigg{)}^{p}\ket{0},$ (52)
@@ B.1 FermiNet
The periodic version of the FermiNet implemented by Cassella et al. [38], which can be found in the FermiNet repository [69], was used as the basis for the AGPs code. The small modifications required to support AGPs wave functions were made using the JAX Python library [46]. For the majority of our calculations, four NVIDIA A100 GPUs were used. For systems with $N>30$ particles, we used four nodes with a total of sixteen A100 GPUs to speed up the calculations. A JAX implementation of the Kronecker-factored approximate curvature (KFAC) gradient descent algorithm [45, 70] was used for optimization. The FermiNet hyperparameters are shown in Table (2) and the network sizes in Table (3). All training runs used $3\times 10^{5}$ iterations to ensure convergence. When evaluating expectation values with an optimized wave function, $5\times 10^{4}$ inference steps were used.
@@
Parameter Value
One-electron Stream Network Size 512
Two-electron Stream Network Size 64
Number of Network Layers 4
Number of Determinants 32
Table 3: Network sizes and number of determinants used in all simulations.

### v4

@@ VI Discussion
In this work, we used neural wave functions to study the benchmark superfluid system known as the UFG 111111 Shortly after submitting this manuscript to the arXiv server, a closely related preprint on neural networks and the Fermi gas appeared [76], proposing a similar extension to continuous-space neural network Ansätzes. . We showed that the Slater FermiNet Ansatz has difficulties in describing paired systems with strong, short-ranged attractive interactions between particles of opposite spin. Hence, we proposed a way to improve the variational Ansatz by using determinants of FermiNet geminals, similar to an AGPs or a BCS wave function. We showed mathematically that the Slater FermiNet is a limiting case of the AGPs FermiNet despite the use of fewer parameters in the latter. It follows that any FermiNet wave function can in principle be written as an AGPs FermiNet wave function.
@@
We compared the total energies and energies per particle of the UFG as calculated using the Slater FermiNet and the AGPs FermiNet. The former fails to produce a paired state when the number of particles, $N$, is greater than around $10$, while the AGPs FermiNet works very well.
@@
As the UFG has a superfluid ground-state, we computed the pairing gap and condensate fraction for the $N=38$ system and compared estimates made with the Slater FermiNet and the AGPs FermiNet. There is a clear qualitative difference between the pairing gap obtained using the AGPs FermiNet and the Slater FermiNet, with the latter approaching zero as the number of particles $N$ increases. Calculations of the superfluid condensate fraction show a similar behavior: the AGPs FermiNet gives an accurate finite result, while the value obtained using the Slater FermiNet tends to zero in the limit of large system size. Although the AGPs pairing gap shows significant finite size errors, it lies close to the mean-field BCS result with Gorkov-Melik-Barkhudarov corrections [61]. Taken together, these results show that the Slater FermiNet is unable to represent large systems with superfluid ground states. The AGPs FermiNet is much more suitable for studying paired systems such as the UFG.
@@
To demonstrate the success of the AGPs FermiNet, we also compared our calculated total energies with state-of-the-art fixed-node diffusion QMC energies obtained using a Jastrow-BCS Ansatz [21]. For all systems with more than a few particles, the AGPs FermiNet achieves lower (i.e., better) variational energies than FN-DMC using the same model interaction and system parameters.
@@
The inability of the Slater FermiNet Ansatz to accurately describe the UFG ground state comes as a surprise because the original FermiNet paper [1] argued that any many-body fermionic wave function can be represented as a single determinant of FermiNet orbitals. However, the mathematical argument relies on the construction of FermiNet orbitals with unphysical discontinuities. Whether or not any wave function can be represented as a single determinant of FermiNet orbitals of the type used in practice, which are differentiable everywhere except at electron-electron and electron-nuclear coalescence points, remains an open question. Another limitation is that the architecture of the FermiNet neural network, which is rather simple, may not be able to represent an arbitrary many-electron FermiNet orbital. Even if a single-determinant Slater FermiNet wave function is general in principle, there is no guarantee that it is equally easy to represent all wave functions. It may be that producing an accurate representation of a paired wave function requires the width and number of layers in the neural network to increase rapidly with system size. Furthermore, if a network of fixed size is used, it may be necessary to increase the number of Slater FermiNet determinants rapidly as the system size increases. The observation that the Slater FermiNet works well when $N\lessapprox 10$ but that the quality of the results degrades rapidly for larger systems, along with the scaling study presented in the final part of the Results section, suggests that this is, in fact, the case.
@@
The AGPs FermiNet introduced in this paper shares many of the strengths of the Slater FermiNet. In particular, there is no need to construct and optimize a new basis set for every new system or particle type. If the AGPs FermiNet proves equally successful in other paired systems, it may now be relatively easy to investigate the importance of pairing in molecules, electron-positron systems, electron-hole liquids, and other $s$-wave superfluids. Another strength of the AGPs FermiNet is the ease with which it is possible to optimize linear combinations of determinants of FermiNet pairing orbitals, such as the one in Eq. (35). This is much more difficult to accomplish with conventional wave functions based on explicit two-electron pairing orbitals or pairing orbitals represented as outer products of single-particle orbitals or basis functions. Just as the many-particle orbitals in a Slater FermiNet radically generalize single-particle orbitals by incorporating electron-electron terms in a permutation-equivariant fashion, so the pairing functions in an AGPs FermiNet generalize BCS-style pairing functions by incorporating the effects of the remaining electrons in a permutation-equivariant fashion.
The AGPs FermiNet introduced here has a straightforward Pfaffian extension and can thus be applied to non-$s$-wave and triplet pairing. Therefore, we expect it to become a powerful tool for understanding strongly correlated non-$s$-wave superfluid and superconducting systems such as Helium-3 or high-$T_{c}$ and $p$-wave superconductors. Finally, our approach is not limited to the FermiNet neural network and can be readily adapted to use more recent architectures such as the Psiformer [71], GLOBE and MOON [72], and DeepErwin [73].
@@ A.1 Fixed-particle Number BCS Ground State
|$\ket{\Psi_{\text{BCS}}}$ $=\prod_{\mathbf{k}}(u_{\mathbf{k}}^{\phantom{\dagger}}+v_{\mathbf{k}}^{\phantom{\dagger}}\hat{c}^{\dagger}_{\mathbf{k},\uparrow}\hat{c}^{\dagger}_{-\mathbf{k},\downarrow})\ket{0}$
$=\bigg{(}\prod_{\mathbf{k}}u_{\mathbf{k}}^{\phantom{\dagger}}\bigg{)}e^{\sum_{\mathbf{k}}\varphi_{\mathbf{k}}^{\phantom{\dagger}}\hat{c}^{\dagger}_{\mathbf{k},\uparrow}\hat{c}^{\dagger}_{-\mathbf{k},\downarrow}}\ket{0},$
@@
|$\ket{\Psi_{\text{PBCS}}}=\bigg{(}\sum_{\mathbf{k}}\varphi_{\mathbf{k}}^{\phantom{\dagger}}\hat{c}^{\dagger}_{\mathbf{k},\uparrow}\hat{c}^{\dagger}_{-\mathbf{k},\downarrow}\bigg{)}^{p}\ket{0},$ (44)
@@ B.1 FermiNet
The periodic version of the FermiNet implemented by Cassella et al. [38], which can be found in the FermiNet repository [77], was used as the basis for the AGPs code. The small modifications required to support AGPs wave functions were made using the JAX Python library [47]. For the majority of our calculations, four NVIDIA A100 GPUs were used. For systems with $N>30$ particles, we used four nodes with a total of sixteen A100 GPUs to speed up the calculations. A JAX implementation of the Kronecker-factored approximate curvature (KFAC) gradient descent algorithm [46, 78] was used for optimization. The initial parameters of the network are initialized using Xavier (random) initialization[79] and the positions of the particles are initialized uniformly in the simulation box. We do not observe significant run-to-run variation in the final energy of a given system as a function of the random initialization. The FermiNet hyperparameters are shown in Table (2) and the network sizes in Table (3). All training runs used $3\times 10^{5}$ iterations to ensure convergence, except for the $N=66$ system (see Appendix F). When evaluating expectation values with an optimized wave function, $5\times 10^{4}$ inference steps were used.
@@
Parameter Symbol Value
One-electron Stream Network Size $n_{l}$ 512
Two-electron Stream Network Size N/A 64
Number of Network Layers $L$ 4
Number of Determinants $D$ 32
Table 3: Network sizes and number of determinants used in all simulations. The corresponding mathematical symbols mentioned in the main text of the paper, where available, are also listed.
@@ B.2 Estimation of the Two-Body Density Matrix
Appendix C How many block-diagonal determinants does the Slater FermiNet need to achieve the same accuracy as the AGPs FermiNet with one determinant?
In the main text, we have demonstrated that the Slater FermiNet with 32 block-diagonal determinants is able to capture superfluidity in small systems but fails at larger systems. Therefore, we are interested in the scaling of the original block-diagonal determinant FermiNet wave function with respect to the system size, i.e. how many block-diagonal determinants do we need in order for the Slater FermiNet to converge to the ground state at each system size?
To answer this question, first we set the AGPs FermiNet with one determinant energies as baselines and plot the percentage difference of the Slater FermiNet from the baseline against the number of block-diagonal determinants used in the Slater FermiNet wave function, repeated at different system sizes from 4 to 14 even particles systems. The results are shown in Fig. 6.
Figure 6: Percentage difference between the Slater FermiNet with various number of block-diagonal determinants and the AGPs FermiNet with one determinant at different system sizes, from 4 to 14 particles.
As the results suggested, it becomes more difficult for the Slater FermiNet to get close to the AGPs FermiNet baseline as the number of determinant increases, especially for larger systems. This is due to the limited performance of the optimizer as the number of determinant increases. Hence, due to the constraints on time and resources, it is not feasible to continue to increase the number of determinants until the Slater FermiNet achieve the same accuracy as the AGPs FermiNet. Instead, we decided to set two thresholds for the percentage difference between the two results. Here, we have used 5% and 10% as the thresholds, plotted as the two horizontal lines in Fig. 6. By plotting the x-intercept of the curves with the two threshold lines against system size, we can determine the relationship between the two, as shown in Fig. 5. As the y-axis in Fig. 5 is in logarithmic scale, a roughly linear relationship suggests an exponential scaling of the number of block-diagonal determinants as system size increases.
The result indicates that, in theory, the Slater FermiNet is capable of converging to the ground state given the number of block-diagonal determinants is sufficiently large. In practice, the number of block-diagonal determinants required to learn the ground state increase exponentially as system size gets bigger and the number get inaccessible rapidly.
Appendix D Dense determinant
Previous works [38, 32, 71, 80, 81, 82] have suggested that the use of dense determinants,
$\Psi^{\text{Dense}}_{\text{FermiNet}}=\det\begin{pmatrix}\phi_{i}^{\uparrow}(\mathbf{r}_{j}^{\uparrow};\{\mathbf{r}_{/j}^{\uparrow}\};\{\mathbf{r}^{\downarrow}\})&\phi_{i}^{\uparrow}(\mathbf{r}_{j}^{\downarrow};\{\mathbf{r}_{/j}^{\downarrow}\};\{\mathbf{r}^{\uparrow}\})\\ \phi_{i}^{\downarrow}(\mathbf{r}_{j}^{\uparrow};\{\mathbf{r}_{/j}^{\uparrow}\};\{\mathbf{r}^{\downarrow}\})&\phi_{i}^{\downarrow}(\mathbf{r}_{j}^{\downarrow};\{\mathbf{r}_{/j}^{\downarrow}\};\{\mathbf{r}^{\uparrow}\})\\ \end{pmatrix}$ (51)
in contrast to block-diagonal diagonal determinants with FermiNet in Eq. 11, provide a slight gain in accuracy in various infinite and periodic systems. In this subsection, we present a small set of calculations of the UFG with dense determinants and compare the results with the block-diagonal FermiNet and the AGPs FermiNet. The results are presented in Table 4.
Type of Total Energy [$E_{FG}$] Pairing
Wave Function $N=36$ $N=37$ $N=38$ Gap [$E_{FG}$]
Block 17.8523(5) 18.2032(5) 18.3131(6) 0.121(2)
Dense 17.7027(5) 18.1601(5) 18.2570(5) 0.180(2)
AGPs 14.6059(4) 15.8060(5) 15.3975(5) 0.804(1)
Table 4: Total energy of the UFG with 36 to 38 particles and their corresponding pairing gap using different wave functions.
Although the use of dense determinants in the UFG does provide lower energies comparing to the block-diagonal determinants, they are still significantly higher than the AGPs FermiNet energies. In addition, the qualitative behaviors, such as the absence of odd-even staggering, are still similar to the block-diagonal FermiNet, which are qualitatively different from the AGPs FermiNet.
Appendix E Translational symmetry of the AGPs FermiNet
In this section, we discuss the consequences of setting an origin in the simulation cell.
As mentioned in the main text, in all of our simulations, an origin is set at one corner of the simulation cell. The particle coordinates $\mathbf{r}^{\alpha}_{i}$, which are used as inputs to the one-electron stream network, are taken with respect to this origin. Thus, the AGPs FermiNet Ansatz does not impose translational invariance with respect to simultaneous translations of all particle coordinates.
As the model unitary Fermi gas system has a translationally invariant ground-state, it is natural to use a translationally invariant Ansatz to study it. Whilst it is possible to neglect the one-electron stream network and only use the two-electron stream network, which is translationally invariant by construction, in the AGPs FermiNet simulations, we found that doing so gives significantly worse energies than the standard AGPs FermiNet. This is due to the limitation on the expressivity of the two-electron stream network, which, despite being translationally invariant, is not sufficient to represent the ground-state of the unitary Fermi gas. This finding is also consistent with the findings in the previous study of the homogeneous electron gas by G. Cassella, et al. [38]. Hence, we include the one-electron stream networks in our Ansatz when studying the unitary Fermi gas.
Figure 7: 2D projection of the $N=14$ unitary Fermi gas density in the simulation cell from sampling the converged AGPs FermiNet wave function. $L$ is the length of the simulation cell.
Despite an explicit origin being embedded into the one-electron stream network in the AGPs FermiNet Ansatz when studying the unitary Fermi gas, we have concluded that this does not create any obvious bias in the converged wave function. Evidence supporting this is shown in Fig. (7), where a 2D-projected density of the $N=14$ unitary Fermi gas, obtained using the converged AGPs FermiNet wave function, is plotted. There is no obvious structure in the density plot, and is uniform across the whole simulation cell. Hence, we believe the Ansatz is capable of representing translationally invariant wave functions, even though this is not explicitly imposed.
Appendix F Training curves of the 66 particle UFG
We present the training curves of the $N=66$ UFG in this section. Due to limited computational resources, we only trained the 66 particle UFG system up to 150,000 steps, which is half the number of steps we use to train all other systems in the paper (see Tab. 2). Therefore, even though our $N=66$ results are able to outperform the FN-DMC results by a similar amount as the smaller systems, it is not fully converged as demonstrated in the training curve below. We emphasize that the energy we obtained is still a variational upper bound on the ground state energy.
Figure 8: Training curve of the 66 particle UFG with 32 determinants and using the geminals in Eq. 19. Red dashed line is the FN-DMC result from [21] under the same density and interaction width.